OpenStudy (anonymous):

Solve (3x)^(log(3))=(4x)^(log(4))

OpenStudy (anonymous):

Log as log base 10

OpenStudy (anonymous):

enjoy, friend. hihi

OpenStudy (anonymous):

\[3^{\log 3}x^{\log 3}=4^{\log 4}x^{\log 4}\\ \frac{x^{\log 3}}{x^{\log 4}}=\frac{4^{\log 4}}{3^{\log 3}}\\ x^{\log 3-\log 4}=\frac{4^{\log 4}}{3^{\log 3}}\\ (\log 3-\log 4)\log x=\log(4^{\log 4})-\log(3^{\log 3})\\ \log x=\frac{(\log 4)^2-(\log 3)^2}{\log 3-\log 4}\\ \log x=\frac{(\log 4-\log 3)(\log 4+\log 3)}{\log 3-\log 4}\\ \log x=-(\log 4+\log 3)=-\log 12=\log {12^-1}\\ x = 12^{-1}\implies \boxed{x={1\over 12}} \]

OpenStudy (anonymous):

@Hoa I did thoroughly... :-)

OpenStudy (anonymous):

amazing!!! how patient you are. hand off

OpenStudy (anonymous):

and yet no medals!!!!!